OpenStudy (anonymous):

125-t^6 ------------- factor it

OpenStudy (anonymous):

\[x^3-y^3=(x-y)(x^2+xy+y^2)\]

OpenStudy (anonymous):

so wuts the answer??

OpenStudy (anonymous):

\[125=(5)^3\] \[t^6=(t^2)^3\]

OpenStudy (anonymous):

so \[(5)^3-(t^2)^3=?\]

OpenStudy (anonymous):

oh ok thnxs

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

but wait thats not factored

OpenStudy (anonymous):

+u need to substitute into the identity that math gave in his first post

OpenStudy (anonymous):

x=5 and y = t^2

OpenStudy (anonymous):

so it would b ( 5-t^2) ( t^4+5t^2+t^4)

OpenStudy (anonymous):

gj :)

OpenStudy (anonymous):

kk and i got another question